OpenStudy (anonymous):

Which is equivalent to Square root of 3 • Square root of -3

OpenStudy (anonymous):

are there choices?

OpenStudy (anonymous):

f 9i g 3i h 9 j -3i

OpenStudy (anonymous):

\[\sqrt{3}\sqrt{-3}=\sqrt{3}\sqrt{3}\sqrt{-1}=3i\]